### v1

# Sparse but Wrong: Incorrect L0 Leads to Incorrect Features in Sparse Autoencoders

Abstract
Sparse Autoencoders (SAEs) extract features from LLM internal activations, meant to correspond to single concepts. A core SAE training hyperparameter is L0: how many features should fire per token on average. Existing work compares SAE algorithms using sparsity–reconstruction tradeoff plots, implying L0 is a free parameter with no single correct value. In this work we study the effect of L0 on BatchTopK SAEs, and show that if L0 is not set precisely, the SAE fails to learn the underlying features of the LLM. If L0 is too low, the SAE will mix correlated features to improve reconstruction. If L0 is too high, the SAE finds degenerate solutions that also mix features. Further, we demonstrate a method to determine the correct L0 value for an SAE on a given training distribution, which finds the true L0 in toy models and coincides with peak sparse probing performance in LLMs. We find that most commonly used SAEs have an L0 that is too low. Our work shows that, to train SAEs with correct features, practitioners must set L0 correctly.
1 Introduction
Sparse autoencoders (SAEs) decompose the dense, polysemantic activations of LLMs into interpretable latent features (Cunningham et al., 2024; Bricken et al., 2023) using sparse dictionary learning (Olshausen and Field, 1997). SAEs have the advantage of being unsupervised, and can be scaled to millions of neurons in its hidden layer (hereafter called “latents”) (Templeton et al., 2024; Gao et al., 2024). When training an SAE, practitioners must decide on the L0 of the SAE: that is, the sparsity; how many latents activate on average for a given input. [^1] TopK and BatchTopK SAEs (Gao et al., 2024; Bussmann et al., 2024) set the L0 ($K$) directly, whereas L1 and JumpReLU (Cunningham et al., 2024; Bricken et al., 2023; Rajamanoharan et al., 2024) adjust it via a coefficient in the loss. In any case, all SAE trainers must decide on the target L0. L0 is typically considered a neutral design choice: most of the literature evaluates SAEs at a range of L0 values, referring to this as a “sparsity–reconstruction tradeoff” (Gao et al., 2024; Rajamanoharan et al., 2024), implying any L0 is equally valid.
However, recent work shows the same trend: too low an L0 leads to worse SAE performance on downstream tasks Kantamneni et al. (2025); Bussmann et al. (2025). In this work, we explore the effect of L0 on SAEs. Starting by exploring a toy model with correlated features, we demonstrate that if the L0 is too low, the SAE can “cheat” by mixing together components of correlated features, achieving better reconstruction compared to an SAE with correct, disentangled features. We consider this to be a manifestation of feature hedging (Chanin et al., 2025), where the SAE abuses feature correlations to compensate for insufficient resources to correctly model the underlying features. Furthermore, we show that if the L0 is too high, the SAE can learn degenerate mixtures of features. We show that it is possible to determine the correct L0 of an SAE by observing projection magnitude between the SAE decoder and training activations.
We validate these findings on Gemma-2-2b (Team et al., 2024), demonstrating that the same decoder patterns we observe in our toy model experiments also manifest in LLM SAEs. We further validate that the optimal L0 we find with our method in Gemma-2-2b matches with peak performance on sparse probing tasks (Kantamneni et al., 2025). Our findings show that L0 must be set correctly for SAEs to learn correct features, and implies that most SAEs used by researchers today have too low an L0.
2 Background
Sparse autoencoders (SAEs).
An SAE decomposes an input activation $\mathbf{x}\in\mathbb{R}^{d}$ into a hidden state $\mathbf{a}$ consisting of $h$ hidden neurons, called “latents”. An SAE is composed of an encoder $\mathbf{W}_{\text{enc}}\in\mathbb{R}^{h\times d}$, a decoder $\mathbf{W}_{\text{dec}}\in\mathbb{R}^{d\times h}$, a decoder bias $\mathbf{b}_{\text{dec}}\in\mathbb{R}^{d}$, and encoder bias $\mathbf{b}_{\text{enc}}\in\mathbb{R}^{h}$, and a nonlinearity $\sigma$, typically ReLU or a variant like JumpReLU (Rajamanoharan et al., 2024), TopK (Gao et al., 2024) or BatchTopK (Bussmann et al., 2024).
$\mathbf{a}=$ $\sigma(\mathbf{W}_{\text{enc}}(\mathbf{x}-\mathbf{b}_{\text{dec}})+\mathbf{b}_{\text{enc}})$ (1)
$\hat{\mathbf{x}}$= $\mathbf{W}_{\text{dec}}\mathbf{a}+\mathbf{b}_{\text{dec}}$ (2)
3 Toy model experiments
We set up a toy model with 50 mutually orthogonal true features $F=\{f_{0},\dots,f_{49}\}\in\mathbb{R}^{100}$. Each feature $f_{i}$ has firing probability $P_{i}$. We set $P(f_{0})=0.345$, and we linearly decrease $P_{i}$ to $P_{49}=0.05$, so feature firing probability decreases with feature number. We randomly generate a correlation matrix, so the firings of each feature are correlated with other features. Features fire with magnitude $\sim\mathcal{N}(1.0,0.15)$. We sampling true feature firings, and summing all firing features to create a training input for the SAE. The goal of the SAE, then, is to learn these true features despite only being trained on their summed activations. Since we have ground-truth knowledge, we know the true number of features firing on average. We call this the true L0, which is 11 for this toy model.
Throughout this work, we use BatchTopK SAEs (Bussmann et al., 2024), as this architecture allows directly controlling L0 rather than controlling it indirectly via a L1 coefficient (Cunningham et al., 2024; Bricken et al., 2023) or L0 coefficient (Rajamanoharan et al., 2024) in the loss. For these toy model experiments, we train SAEs on 15M synthetic samples using SAELens (Bloom et al., 2024).
(a) SAE L0 = 5
(b) SAE L0 = 11
(c) SAE L0 = 18
Figure 1: SAE decoder cosine similarity with true features. True features L0 = 11.
We train SAEs with L0 that is too low (L0=5), exactly correct (L0=11), and too high (L0=18). Results are shown in Figure 1. When the SAE L0 matches the true L0 (Figure 1(b)), the SAE exactly learns the true features. When the SAE L0 is too low (Figure 1(a)), the SAE mixes components of correlated features together, especially breaking latents tracking high-frequency features. When SAE L0 is too high (Figure 1(c)), the SAE learns degenerate solutions that mix features together. The further the L0 is from the true L0, the more broken the SAE becomes. Interestingly, when L0 is too high the SAE still learns many correct latents, but when L0 is too low, every latent in the SAE is affected.
3.1 MSE loss incentivizes low-L0 SAEs to learn incorrect features
Why do SAEs with too low L0 not learn the true features? We take the correct SAE from Figure 1(b) and set L0=5, to match the SAE from Figure 1(a). We then generate 100k training samples and calculate the Mean Square Error (MSE) of both these SAEs. The SAE with incorrect latents from Figure 1(a) achieves a MSE of $2.73$, while the SAE with ground-truth correct latents achieves a much worse MSE of $4.88$. Thus, MSE loss actively incentivizes low L0 SAEs to learn incorrect latents. Furthermore, this shows that using reconstruction error as a target, and thus the sparsity–reconstruction tradeoff (Gao et al., 2024; Rajamanoharan et al., 2024; Cunningham et al., 2024) is harmful: when the L0 is too low, the incorrect SAE achieves better reconstruction than the ground-truth correct SAE. We explore this further in Appendix A.2.
3.2 Detecting the true L0 using the SAE decoder
Figure 1 reveals that the SAE decoder latents contain mixes of underlying features, both when the L0 is too high and also when it is too low. As the SAE approaches the correct L0, each SAE latent has fewer components of multiple true features mixed in, becoming more monosemantic. Thus, we expect that when the SAE has the correct L0, most latents should have near zero projection on arbitrary training inputs, because they usually do not contain the feature being tracked by that latent. If we are far from the correct L0, then SAE latents contain components of many underlying features, and we expect latents to project unexpectedly strongly on arbitrary training inputs.
We now define a metric we call Nth decoder projection score, or $s_{n}^{\text{dec}}$, that we can minimize to find the optimal L0 of the SAE. Given SAE inputs $\mathbf{x}\in\mathbb{R}^{b\times d}$ where $b$ is the batch size and $d$ is the input dimension, we first compute the decoder projections for all latents:
$\mathbf{Z}=(\mathbf{x}-\mathbf{b}_{\text{dec}})\mathbf{W}_{\text{dec}}^{\top}\in\mathbb{R}^{b\times h}$ (3)
where $\mathbf{b}_{\text{dec}}\in\mathbb{R}^{d}$ is the decoder bias and $\mathbf{W}_{\text{dec}}\in\mathbb{R}^{d\times h}$ is the decoder weight matrix with $h$ latent dimensions. To aggregate across the batch, we flatten $\mathbf{Z}$ to obtain $\mathbf{z}\in\mathbb{R}^{bh}$ and sort these values in descending order to get $\mathbf{z}_{\downarrow}$. The Nth decoder projection is then defined as:
$s_{n}^{\text{dec}}=\mathbf{z}_{\downarrow}[n\cdot b]$ (4)
where $n\in\{1,2,\ldots,h\}$ indexes the latent by its ranking. The multiplication by $b$ accounts for the batch dimension, effectively selecting the $n$th highest projection value when considering all samples in the batch. For this to work $n$ should be sufficiently larger than a reasonable guess at the correct L0, as in a perfect SAE, the decoder for these latents should be uncorrelated with input activations. Empirically, picking $n$ up to about $\nicefrac{{h}}{{2}}$ seems to work well.
Figure 2: Nth decoder projection vs SAE L0 for N=12 (left) and N=20 (right) on our toy model SAEs. The true L0, 11, is marked by a dotted line on the plots. Both settings of N are minimized at the true L0, although the slopes of the metric are slightly different depending on N.
We calculate $s_{n}^{\text{dec}}$ for $n=18$ and $n=22$, with SAE L0 ranging from 2 to 25 in Figure 2. We see that the metric is minimized at the true L0, 11, in both cases, although the slope of the metric changes depending on $n$. In both cases, the slope of $s_{n}^{\text{dec}}$ is flat when L0 is slightly higher than the true L0.
4 LLM experiments
Next, we train a series of BatchTopK SAEs (Bussmann et al., 2024) on Gemma-2-2b (Team et al., 2024) layer 12 with different L0 values and calculate $s_{n}^{\text{dec}}$ for these SAEs. Each SAE is trained on 500M tokens from the Pile (Gao et al., 2020) using SAELens (Bloom et al., 2024) with a learning rate of $3e^{-4}$. During training, we ensure that the decoder remains normalized with $W_{dec}||_{2}=1$ so $s_{n}^{\text{dec}}$ calculations use the same scale for every latent. We train an SAE at each L0 from $10,20,30,\dots,250$. After L0=250, due to compute costs, we increase the spacing of the L0 of the SAEs up to L0=2500.
Figure 3: Nth decoder projection vs SAE L0 for N=700 (left) and N=10k (right) on Gemma-2-2b layer 12 SAEs with 32k width. Here as well, we see that $s_{n}^{\text{dec}}$ is minimized at the same point regardless of the choice of N, but the slope and shape of the $s_{n}^{\text{dec}}$ curve varies depending on the choice of N. These plots imply that the correct L0 for these SAEs is around 200-250, much higher than typically used.
We plot $s_{n}^{\text{dec}}$ vs L0 for Gemma-2-2b SAEs in Figure 3 for N=700 and N=10k (more plots in Appendix A.7). We see the same pattern for these SAEs as we saw in our toy model SAEs, namely that the metric is minimized at roughly the same L0, 200-250, despite the choice of N. As in the toy model SAEs case, using a higher N value results in a shallower slope leading up the the minimum $s_{n}^{\text{dec}}$ L0.
We test these SAEs using the sparse probing benchmark from Kantamneni et al. (2025), consisting of over 100 sparse probing tasks. We use the normal setting from this benchmark with k=1 and k=16 sparse probes. Results are shown in Figure 4. The results closely match what the $s_{n}^{\text{dec}}$ metric predicts, showing that L0 around 200-250 yields the highest results on the sparse probing benchmark.
Figure 4: Sparse probing F1 score vs L0 for Gemma-2-2b layer 12 SAEs. Results are shown for k=1 (left) and k=16 (right). Both results are maximized at L0 around 200-250, matching our $s_{n}^{\text{dec}}$ findings.
5 Related work
Chanin et al. (2025) explores feature hedging, showing SAEs mix correlated features into latents if the SAE is too narrow. We consider our work a version of feature hedging due to low L0. Till (2024) shows SAEs may increase sparsity by inventing features. Chanin et al. (2024) discuss the problem of feature absorption, where SAEs can improve their sparsity score by mixing hierarchical features together. Engels et al. (2024) investigates SAE errors and finds that SAE error may be pathological and non-linear.
6 Discussion
While most practitioners of SAEs understand that having too high of a L0 is problematic (at some point the sparse autoencoder is not sparse), our work shows that having too low of L0 is perhaps even worse. Our work has several important implications for the field. First, the L0 used by most SAEs is lower than it ideally should be, as a cursory search of open source SAEs on Neuronpedia (Lin, 2023) shows L0 less than 100 and even less than 50 is very common even for SAEs trained on large models (see Appendix A.5). Our work further shows that the framing of a sparsity–reconstruction tradeoff, as commonly discussed by most SAE papers (Cunningham et al., 2024; Gao et al., 2024; Rajamanoharan et al., 2024), is a misleading metric: when L0 is too low, an SAE with a correct dictionary achieves worse reconstruction than an incorrect SAE that mixes components of correlated features (see Appendix A.2).
While we expect that the correct L0 for a given SAE depends on the dataset used to train the SAE, we are also excited that our analysis may give us hints as to the true number of firing features in superposition in any given LLM activation.
The next step in our work is use our $s_{n}^{\text{dec}}$ metric to located the correct L0 for an SAE during training. We discuss one such technique in Appendix A.4. We also hope to estimate how the correct L0 varies with SAE and model size and layer, and investigate more SAE architectures.
Acknowledgments and Disclosure of Funding
David Chanin was supported thanks to EPSRC EP/S021566/1 and the Machine Learning Alignment and Theory Scholars (MATS) program.
Appendix A Technical Appendices and Supplementary Material
A.1 Toy model details
Figure 5: (left) random correlation matrix and (right) base feature firing probabilities for toy model.
A.2 Sparsity–reconstruction tradeoff in toy models
It is common in previous SAE work to discuss a sparsity–reconstruction tradeoff[Cunningham et al., 2024, Gao et al., 2024, Rajamanoharan et al., 2024], where the assumption is that having better reconstruction at a given sparsity is inherently better, and is a sign the SAE is correct. As we discussed in Section 3.1, when the L0 of the SAE is lower than optimal, the SAE can find ways to “cheat” and get a better MSE score by mixing components of correlated features together. This results in an SAE where the latents are no longer monosemantic, and do not track ground-truth features.
Figure 6: Sparsity ($x$) vs reconstruction ($y$) for learned SAEs and a ground-truth correct SAE at a given L0 in our toy model. When L0 is lower than the true L0 of the toy model, the learned SAE gets a higher reconstruction score than the ground-truth correct SAE, highlighting the problems with optimizing for reconstruction as a proxy for the quality of the SAE.
We explore the sparsity–reconstruction tradeoff in SAEs trained on our toy model at various L0s. Since we have the ground-truth correct features in our toy model, we also construct a ground-truth SAE that perfectly represents these true features. We manually set the L0 of the ground truth SAE while leaving the encoder and decoder fixed at the correct features. We plot the variance explained vs L0 in Figure 6 for the learned and ground-truth SAEs. We see that when the L0 of SAE is lower than the true L0 of the toy model, the ground-truth correct SAE actually scores worse on reconstruction than the incorrect learned SAE! This highlights the problems with optimizing reconstruction as a proxy for SAE quality.
(a) Learned SAE L0 = 1
(b) Learned SAE L0 = 2
(c) Ground truth SAE
Figure 7: SAE decoder cosine similarity with true features for the learned SAEs with L0=1 and L0=2, compared with the ground-truth correct SAE. The learned SAEs score more than twice as a high as the ground truth SAE on variance explained when L0=1 and L0=2, despite having extremely corrupted, polysemantic latents.
We show the cosine similarity of the SAE decoder latents with the ground truth features for the SAEs learned with L0=1 and L0=2 compared with the ground-truth SAE in Figure 7(c). Both these SAEs outperform the ground-truth SAE on variance explained by over 2x despite learning horribly polysemantic latents bearing little resemblance to the underlying true features of the model.
A.3 Transitioning L0 during training
We explore the effect of transitioning the L0 of the SAE during training using the toy model from Section 3. This toy model has a true L0 of 11. We train BatchTopK SAEs with a final L0 of 11, but starting with L0 either too high or too low, and linearly transitioning to the correct L0 over the first 25k steps of training, leaving the SAE at the correct L0 for the final 5k steps of training. We use a starting L0 of 20 for the case where we start too high, and use a starting L0 of 2 for the case where we start too low. Results are shown in Figure 8.
Figure 8: Transitioning L0 from too low (left) and too high (right) to the correct L0 during training. When the starting L0 is too high, the SAE still learns the correct features at the end of training. However, when L0 is too low, the SAE cannot recover fully and still learns many incorrect features at the end of training.
We see that decreasing the L0 of the SAE from a too high value to the correct value still results in the SAE learning correct features. However, when the SAE starts from a too low L0, the SAE cannot fully recover when the L0 is adjusted to the correct value later. It seems that the latents the SAE learns when L0 is too low is a local minimum that is difficult from the SAE to escape from even when the L0 is later corrected. This is likely because the latents learned when L0 is too low are optimized by gradient descent to achieve a higher MSE loss than is achievable by the correct latents under the same L0 constraint. However, when L0 is too high, there is no equivalent optimization pressure, and is thus less likely to be a local minimum.
A.4 Automatically finding the correct L0 during training
A natural next step of our finding that the correct L0 occurs when Nth decoder projection, $s_{n}^{\text{dec}}$, metric is minimized is to use this to find the correct L0 automatically during training. This is a meta-learning task, as the L0 is a hyperparameter of the training process. We find there are several challenges to directly using $s_{n}^{\text{dec}}$ as an optimization target:
  • Small gradients directly above correct L0 In our plots of $s_{n}^{\text{dec}}$ from both toy models and Gemma-2-2b, we find that the metric is relatively flat in a region start at the correct L0 and extending to higher L0 values. We thus need a way to traverse this flat region and stop once the metric starts to increase again.
  • The impact of changing L0 is delayed We find that it takes many steps after changing L0 for $s_{n}^{\text{dec}}$ to also change, meaning it is easy to overshoot the target L0 or oscillate back and forth.
  • Dropping L0 too low can harm the SAE As we saw in Appendix A.3, if the L0 is too low the SAE can permanently end up in poor local minima. We thus want to avoid dropping below the correct L0, even temporarily, to avoid permanently breaking the SAE. We therefore need to start with L0 too high and slowly decrease it until we find the correct L0.
  • Noise during training We find that while $s_{n}^{\text{dec}}$ shows clear trends after training for many steps, it can be noisy on each training sample. So our optimization needs to be robust to this noise.
Taking these requirements into account, we present an optimization procedure to find the L0 that minimizes $s_{n}^{\text{dec}}$ automatically during training. We first estimate the gradient of $s_{n}^{\text{dec}}$, hereafter referred to as to as the metric, $m$, with respect to L0, $\nicefrac{{dm}}{{dL0}}$. We first define an evaluation step $t$ as a set number of training steps (we evaluate every 100 training steps). At $t$ we change L0 by $\delta_{L_{0}}$. At the next evaluation step, $t+1$, we evaluate $m$. We use a sliding average of $s_{n}^{\text{dec}}$ over the past 10 training steps to calculate $m$ to help account for noise. We the estimate $\nicefrac{{dm}}{{dL0}}$ as:
$\frac{dm}{dL0}=\frac{m_{t+1}-m_{t}}{\delta_{L0}}$
Next, we add a small negative bias to this gradient estimate to encourage our estimate to push L0 lower even if the loss landscape is relatively flat. We use a bias magnitude $0<b<1$ that is multiplied by the magnitude of our gradient estimate, so that our biased estimate can never change the sign of the gradient estimate, but can gently nudge it to be more negative in flat, noisy regions of the loss landscape. We find $b=0.1$ works well. Thus, our biased gradient estimate $\nicefrac{{dm_{b}}}{{dL0}}$ is calculated as below:
$\frac{dm_{b}}{dL0}=\frac{dm}{dL0}-b\left|\frac{dm}{dL0}\right|$
We then provide this gradient to the Adam optimizer [Kingma and Ba, 2014] with default settings, and allow it to change the L0 parameter.
We add the following optional modifications to this algorithm. First, we clip the gradient estimates $\nicefrac{{dm}}{{dL0}}$ to be between -1 and 1. We also set a minimum and maximum $\delta_{L_{0}}$. The minimum is added to avoid the denominator of our gradient estimate being near 0, and the maximum is chosen to keep the L0 from changing too quickly. In practice, we find a minimum $\delta_{L_{0}}$ between 0 and 1 seems to work well, and a maximum $\delta_{L_{0}}$ between 1 and 5 seems to work well.
We find that this optimization strategy works very well in toy models, but requires a lot of hyperparameter tuning to work in real LLMs. The starting L0, $n$ for $s_{n}^{\text{dec}}$, $b$, learning rate for the Adam optimizer, and min and max $\delta_{L_{0}}$ values all have a big impact on how fast and how aggressively the optimization works. The slop of $m$ around the correct L0 is shallow, so it is easy to overshoot. We also find that different values of $n$ take more or less time to converge during training. We expect it is possible to further simplify and improve this process in future work.
A.5 L0 of open-source SAEs on Neuronpedia
We analyze common open-source SAEs as provided by Neuronpedia [Lin, 2023] and SAELens [Bloom et al., 2024]. We include all SAEs cross-listed in both SAELens and Neuronpedia with an L0 reported in SAELens. We show the results as a histogram in Figure 9. Our analysis shows that for layer 12 of Gemma-2-2b, the correct L0 should be around 200-250. However, we find that most open-source SAEs have L0 below 100, much lower than our analysis expects to be ideal.
Figure 9: L0 of SAEs on Neuronpedia with known L0 listed in SAELens.
A.6 Limitations
We limit our investigated to Gemma-2-2b layer 12 due to the cost associated with training so many SAEs at different L0s. We investigate only BatchTopK SAEs, due to these SAEs allowing direct control over the L0 of the SAE without being moderated through a sparsity coefficient as in L1 SAEs [Cunningham et al., 2024, Bricken et al., 2023] or JumpReLU SAEs [Rajamanoharan et al., 2024], but do not expect our findings to be meaningfully different for these other architectures.
A.7 Extended Gemma-2-2b Nth decoder projection plots
Figure 10: Extended Nth decoder projection plots. Gemma-2-2b, layer 12, 32k latents. Regardless of the choice of N, all plots are minimized around the same L0 range, 200-250.